OpenStudy (anonymous):

Which sum or difference identity would you use to verify that cos (180° - q) = -cos q?

OpenStudy (whpalmer4):

How about \[\cos(a-b) = \cos a \cos b + \sin a \sin b\] with a = 180 degrees?

OpenStudy (anonymous):

It's not even worth trying to explain it cx Just tell me? cx

OpenStudy (anonymous):

Opps didn't realize that was the answer >_<

OpenStudy (whpalmer4):

I just looked down a list of sum and difference identities until I saw one that had cos (a-b)...what's there to explain? :-)

OpenStudy (anonymous):

>_< wow I'm dumb, and idk

OpenStudy (whpalmer4):

What you might need is a handy trig reference sheet. Here's one: http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

thanks c: